How long will it take for $8200 to grow to $46,400 at an interest rate for 10.5% if the interest is compounded continuously?
Answer by lwsshak3(11628)

How long will it take for $8200 to grow to $46,400 at an interest rate for 10.5% if the interest is compounded continuously?
**
Interest formula for continuous compounding: A=Pe^rt, P=initial investment, r=annual interest rate, A=amount after t years.
46400=8200e^.105t
464/82=e^.105t
take log (base e) of both sides using calculator
ln(464/82)=.105t*lne
lne=log of base=1
t=ln(464/82)/.105
t≈16.5 years
ans:For $8200 to grow to $46,400 will take 16.5 years